gnment [ECN221] Exercise 03.33 Algorithmic <Question 8 of 18 Check My Work (2 remaining) eBook Scores turned in by an amateurs Exercise 03.35 Question 9 of 18 Check My Work (2 remaining) eBook Consider a sample with data values of 10, 20, 12, 17, andI need help with both questions
Answer #1

Date meun. of 2014. Exi 73 +77 +78 +76+74+72+74+76 8. = 600 x = 75 mean = 75 standard devinzon xix (xi-D2 73 77 78 76 74 72 7. 2014 mean 75 standard deration 11.94 mean of 2015. Iyi - 70 +69+74+76 +84 +79+70+ 78 - Y= TS i mean= 75] 5 25 36 74 -1 79 o8= {lyi-7² I 2 = 4.92 : 2015 mean 175 steinanal devintion = 4.92 (b) primory differnce is variability is increases Scores areExcersise 03. 35 Data volue 10,20,12, 17 and 16 we have to find first ML 2. .: u and 2 M = 10 +20 +1 2+17+16 5 - 75 es = cхirz score formula = x-4 observed rolne X-10 ..-x-il - 10-15 3.58 = -1.40 X = 2o z = +1.40 = x-el 6 2015 3.58 x = 12 z = 12-15 -Date observed value Z-slose T-1. 40 1.40 -0.84 0.84 0.28